# Applying statistical learning theory to deep learning

### v1

@@ 5.6 Beyond linear models
where $\nabla_{h}L(h)$ is now an element of $\mathbb{R}^{\mathcal{X}}$ and $\nabla F^{\top}\nabla F$ is a linear map from $\mathbb{R}^{\mathcal{X}}:\mathbb{R}^{\mathcal{X}}$, with a kernel taking the form
@@ 6 Lecture 6: Implicit bias with linear functionals and the logistic loss
Summary of the previous lecture
We’re trying to understand how the choice of optimization geometry (i.e., what our preferred metric is in local updates) affects where the optimization will lead us. Last time, we also spoke about geometry of parameters space (usually just Euclidean geometry), but really what mattered was the geometry in function space. The relationship between parameter and function space is:
$h_{w}(x)=f(w,x)\,\text{ or }h_{w}(x)=\langle\beta_{w},x\rangle\,\text{where }\beta_{w}=F(w)\,.$
@@
$\dot{\beta}=-\nabla F^{T}\nabla F\nabla_{\beta}L_{S}(\beta)\,.$
@@
To finish our recap from last lecture, the simplest example in which we can see non-trivial behavior is this squared parameterization model: $f(w,x)=\langle\beta_{w},x\rangle$ with $\beta_{w}=F(w)=w_{+}^{2}-w_{-}^{2}\,,$ or in the deeper case: $\beta_{w}=F(w)=|w_{+}|^{D}-|w_{-}|^{D}\,.$ We talked about initializing at $w(0)=\alpha\vec{1}\,,$ which gives $\beta(0)=0\,.$ We saw that for $D=2\,,$ for large $\alpha\,,$ we got the kernel regime, and when we take $\alpha$ to 0, we get this $\ell_{1}$ regularization. Even when $D\geq 2\,,$ when we took $\alpha\rightarrow 0\,,$ we still get $\ell_{1}$ regularization, despite explicit $\ell_{2}$ norm regularization on the parameters not giving $\ell_{1}$ norm regularization in function space.
Today, we will look at logistic loss and see that there is an effect in the classification setting, as well.
@@ 6.1 Problem setting and equivalent reformulation
$\mathbf{w}_{k+1}=\mathbf{w}_{k}-\eta\nabla\mathcal{L}_{S}(\mathbf{w}_{k})$ (138)
In the overparametrized case, i.e. $d>n$, the data is separable and we may minimize the loss by considering any separating hyperplane and taking its norm to infinity. Since the optimal solution diverges, we focus on the direction of the optimal solution :
@@
The Lagrangian for this problem reads
@@ 6.2 Gradient flow dynamics
$\mbox{log}(1+\mbox{exp}(-\hat{y}_{i}y_{i}))\simeq\mbox{exp}^{-y_{i}\hat{y}_{i}}\quad\forall i$ (144)
The gradient reads
$-\nabla_{{\mathbf{w}}}L({\mathbf{w}})=\frac{1}{n}\sum_{i=1}^{n}e^{-\langle{\mathbf{w}}_{i},\boldsymbol{x}_{i}\rangle}\boldsymbol{x}_{i}$ (145)
Formal proof of what follows can be found in [18, 21]. Intuitively, GF finds a separating direction that will not change much after a certain number of iteration, and increase its norm. We may then write
@@
where $\rho(k)=o(1)$ (a Theorem from [Soudry Hoffer Srebro 18] actually shows $g(k)=log(k)$). Replacing in the expression of the gradient leads to
$\nabla_{{\mathbf{w}}}L({\mathbf{w}})=\frac{1}{n}\sum_{i=1}^{n}e^{-\langle{\mathbf{w}},\boldsymbol{x}_{i}\rangle}\boldsymbol{x}_{i}\simeq\frac{1}{n}\sum_{i=1}^{n}e^{g(k)\langle{\mathbf{w}}_{\infty},\boldsymbol{x}_{i}\rangle-O(1)}\boldsymbol{x}_{i}$ (147)
@@
whose norm will remain finite. Primal feasibility is verified by construction $\langle{\mathbf{w}},\boldsymbol{x}_{i}\rangle\geqslant 1$ is satisfied by construction, and we have seen that the zero gradient condition on ${\mathbf{w}}$ prescribing it as a linear combination of the data points is also verified. We also see that large values of $\langle{\mathbf{w}}_{\infty},\boldsymbol{x}_{i}\rangle$ the corresponding coefficient $\nu_{i}$ will decrease very fast as $g(k)\to\infty$, leaving the main contribution to the lowest values which are the $\boldsymbol{x}_{i}$ for which $\langle{\mathbf{w}}_{\infty},\boldsymbol{x}_{i}\rangle=1$. Thus we recover the complementary slackness condition.
@@ 6.3 Comparing the squared, logistic and exponential loss
The first order stationary points of this objective are exactly those that satisfy the KKT conditions. For convex problems, this implies optimality. Here, we have to be a bit more careful. The objective is convex but $f$ is not in general a convex function of $w$. This suggests that the implicit bias is defined by $R_{F}(h)=\arg\min_{F(w)=h}\lVert w\rVert_{2}\,,$ i.e., which function $h$ is representable most cheaply in parameter space. These two problems have the same global minimum but this does not mean a first-order stationary point of the first problem are first-order stationary points of the second problem. Relating the two remains largely open.
@@
For squared loss, as the scaling goes to infinity, the entire trajectory converges to the kernel trajectory. In particular, then, the implicit bias is the implicit bias of the kernel (i.e., the limit points are the same). For logistic loss, we can get a similar statement about infinite scale leads to the kernel regime, but the problem is that we can only get it for finite time. That is, if we fix the amount of time for which we optimize and then take the scale to infinity, then we stay inside the kernel regime. These results are presented formally in [8]. For logistic loss, if we change the order of limits for scale of initialization and time of optimization, we reach a first order stationary point of $\arg\min\lVert w\rVert_{2}$ such that the margin condition is satisfied (formal statement in [18] and [22]).
@@
Thus, for logistic loss, the regime we are in is no longer just dependent on $\alpha$, the initialization scale, but rather on the combination of scale and training loss. For each optimization accuracy, we can ask at what scale we would enter the kernel regime. The transition occurs at $\epsilon\sim\exp{(-\alpha^{2})}$. On the boundary, (under several non-realistic assumptions), the behavior follows the $Q_{\alpha}$ function with parameter $\frac{\alpha}{\sqrt{\log(1/\epsilon)}}\,.$ Thus, the transition depends on the ordering of $\epsilon,\alpha$ limits. While it is true that we have an asymptotic result, it only kicks in when $\epsilon=10^{-1700}\,,$ so we do not get $\ell_{1}$ regularization in practice.
@@
Figure 20: This figure depicts how the predictor behaves with finite but not large initialization scale ($\alpha=2$). We know that in the long run, we will reach the minimum $\ell_{1}$ norm solution. When the optimization accuracy is finite, we will traverse the whole $Q_{\alpha}$ path. The dashed black line is the path of minimum $Q_{\alpha}$ margin solution: solution with margin 1 that minimizes $Q_{\alpha}$ for corresponding $\alpha$. As $\alpha$ increases, it seems we converge to this path. (Note, we don’t know how to prove this but it seems to hold empirically.) Importantly, while it is true asymptotically in $\epsilon\,$ that minimizing the regularizer in function space corresponds to minimizing the norm of the weights, the $\epsilon$ at which it starts to hold is completely impractical.
@@
Other control parameters (aside from initialization scale) include how early we stop training, shape of the initialization (i.e., relative scale of the parameters), step size, and stochasticity. The latter has been studied through the lens of batch size and label noise.t
@@ 6.4 Matrix Factorization Setting and Commutativity
We are still going to look at a linear model in $\beta$, but now $\beta$ is going to be a matrix. This is a standard least squares objective in $\beta$. The main difference is that now, instead of factorizing $\beta$ element-wise, we are going to factorize $\beta=UV^{T}$. This formulation captures many things: matrix completion, matrix reconstruction from linear measurements, and even multi-task learning.
@@
Observe that this is a linear transformation of the gradient. We would be interested in writing the transformation in terms of just $\beta$, and this is not possible here, as it turns out. Instead, we define:
@@
The local geometry of the space in which we search is given by left and right multiplying by $\tilde{\beta}\,.$ Now what we want to see if we can identify this as a Hessian map. Do the dynamics stay in a low-dimensional manifold?
As before, where we are depends on the integral of the residual so far. The residual explains how much to multiply by the observation. If matrices commute, we can solve the differential equation explicitly with the matrix exponential. The metric tensor is a Hessian map, which directly implies we are in low-dimension manifold. In this case, the result is very robust: it doesn’t depend on the residuals, nor the loss under which the residuals are computed. It is not robust to step size: could do a ski jump off the manifold now that it is curved. If it is non-commutative, it matters which order in which I multiply on right and left. $\beta(t)\,,$ is a “time-ordered exponential,” and we cannot ignore the ordering of the residuals. Even with just two data points, we can enter the whole space. An analogy is parallel parking, where with the forward/backward and left/right controls we can somehow move net to the right.
@@
7 Lecture Series Summary
The approach covered over this lecture series involved attempting to characterize the implicit bias of neural networks from a complexity theory perspective. The main question is to understand which zero-error solution is reached when optimizing an overparameterized problem, and determining this involved decomposing the problem into (1) identify what optimization biases toward based on identifying what complexity measure is being (at least approximately) minimized; followed by (2) evaluating whether this complexity measure is a good one to optimize for the task at hand, i.e., does a minimizer of it generalize well.
A natural question, then, is whether this is the right approach to take. Indeed, it is not always the best description for optimization in neural networks. It remains open whether minimization of this complexity measure, perhaps even approximately, is sufficient to explain most cases.
The ultimate questions: what is true Inductive Bias? What makes reality efficiently learnable by fitting a (huge) neural net with a specific algorithm?
@@ Acknowledgements
These are notes from the lecture of Nathan Srebro given at the summer school “Statistical Physics & Machine Learning ”, that took place in Les Houches School of Physics in France from 4th to 29th July 2022. The school was organized by Florent Krzakala and Lenka Zdeborová from EPFL.

### v2

@@ 5.6 Beyond linear models
where $\nabla_{h}L(h)$ is now an element of $\mathbb{R}^{\mathcal{X}}$ and $\nabla F^{\top}\nabla F$ is a linear map from $\mathbb{R}^{\mathcal{X}}$ to $\mathbb{R}^{\mathcal{X}}$, with a kernel taking the form
@@ 6 Lecture 6: Implicit bias with linear functionals and the logistic loss
Recall that we are trying to understand how the choice of optimization geometry (i.e., what our preferred metric is in local updates) affects where the optimization will lead us. In the previous lecture, we also spoke about geometry of parameters space (usually just Euclidean geometry), but really what mattered was the geometry in function space. Recall that the relationship between parameter and function space is:
$h_{w}(x)=f(w,x)\,\text{ or }h_{w}(x)=\langle\beta_{w},x\rangle\,$
where $\text{where }\beta_{w}=F(w)\,.$ (134)
@@
$\dot{\beta}=-(\nabla F^{T}\nabla F)\nabla_{\beta}L_{S}(\beta)\,.$
@@
The simplest example in which we can see non-trivial behavior is this squared parameterization model: $f(w,x)=\langle\beta_{w},x\rangle$ with $\beta_{w}=F(w)=w_{+}^{2}-w_{-}^{2}\,,$ or in the deeper case: $\beta_{w}=F(w)=|w_{+}|^{D}-|w_{-}|^{D}\,.$ We talked about initializing at $w(0)=\alpha\vec{1}\,,$ which gives $\beta(0)=0\,.$ We saw that for $D=2\,,$ for large $\alpha\,,$ we got the kernel regime, and when we take $\alpha$ to 0, we get this $\ell_{1}$ regularization. Even when $D\geq 2\,,$ when we took $\alpha\rightarrow 0\,,$ we still get $\ell_{1}$ regularization, despite explicit $\ell_{2}$ norm regularization on the parameters not giving $\ell_{1}$ norm regularization in function space.
In this lecture, we will move away from the least-squares setting and look at logistic regression. We will see that there is an implicit bias of optimization in the classification setting, as well.
@@ 6.1 Problem setting and equivalent reformulation
$\mathbf{w}_{k+1}=\mathbf{w}_{k}-\eta\nabla L_{S}(\mathbf{w}_{k})$ (136)
In the overparametrized case, i.e. $d>n$, the data is separable and we may minimize the loss by considering any separating hyperplane and taking its norm to infinity (the loss does not have a finite minimizer in this case). Since the optimal solution diverges, we focus on the direction of the optimal solution :
@@
The Lagrangian, denoted $\mathcal{L}$, for this problem reads
@@ 6.2 Gradient flow dynamics
$\forall i\quad\mbox{log}(1+\mbox{exp}(-\hat{y}_{i}y_{i}))\simeq\mbox{exp}^{-y_{i}\hat{y}_{i}}$ (142)
The gradient may then be approximated by
$-\nabla_{{\mathbf{w}}}L_{S}({\mathbf{w}})=\frac{1}{n}\sum_{i=1}^{n}e^{-\langle{\mathbf{w}}_{i},\boldsymbol{x}_{i}\rangle}\boldsymbol{x}_{i}.$ (143)
The formal proof of what follows can be found in [32, 35]. Intuitively, GF finds a separating direction that will not change much after a certain number of iteration, and increase its norm. We may then write
@@
where $\rho(k)=o(1)$ (a Theorem from [Soudry Hoffer Srebro 18] actually shows that $g(k)=log(k)$). Replacing in the expression of the gradient leads to
$\nabla_{{\mathbf{w}}}L_{S}({\mathbf{w}})=\frac{1}{n}\sum_{i=1}^{n}e^{-\langle{\mathbf{w}},\boldsymbol{x}_{i}\rangle}\boldsymbol{x}_{i}\simeq\frac{1}{n}\sum_{i=1}^{n}e^{g(k)\langle{\mathbf{w}}_{\infty},\boldsymbol{x}_{i}\rangle-O(1)}\boldsymbol{x}_{i},$ (145)
@@
whose norm will remain finite. Primal feasibility is verified by construction $\langle{\mathbf{w}},\boldsymbol{x}_{i}\rangle\geqslant 1$ is satisfied by construction, and we have seen that the zero gradient condition on ${\mathbf{w}}$ prescribing it as a linear combination of the data points is also verified. We also see that for large values of $\langle{\mathbf{w}}_{\infty},\boldsymbol{x}_{i}\rangle$ the corresponding coefficient $\nu_{i}$ will decrease very fast as $g(k)\to\infty$, leaving the main contribution to the lowest values which are the $\boldsymbol{x}_{i}$ for which $\langle{\mathbf{w}}_{\infty},\boldsymbol{x}_{i}\rangle=1$. Thus we recover the complementary slackness condition.
@@ 6.3 Comparing the squared, logistic and exponential loss
The first order stationary points of this objective are exactly those that satisfy the KKT conditions. For convex problems, this implies optimality. Here, we have to be a bit more careful. The objective is convex but $f$ is not in general a convex function of $w$. This suggests that the implicit bias is defined by $R_{F}(h)=\arg\min_{F(w)=h}\lVert w\rVert_{2}\,,$ i.e., which function $h$ is representable with the smallest possible norm in parameter space. These two problems have the same global minimum but this does not mean that a first-order stationary point of the first problem is a first-order stationary point of the second problem. Relating the two remains largely open.
@@
For squared loss, as the magnitude of initialization goes to infinity, the entire trajectory converges to the kernel trajectory. In particular, then, the implicit bias is the implicit bias of the kernel (i.e., the limit points are the same). For the logistic loss, we can get a similar statement about the infinite scale leading to the kernel regime, but only for finite time. That is, if we fix the amount of time for which we optimize and then take the scale to infinity, then the kernel regime will still be observed. These results are presented formally in [16]. For the logistic loss, if we change the order of limits for scale of initialization and time of optimization, we reach a first order stationary point of $\arg\min\lVert w\rVert_{2}$ such that the margin condition is satisfied (formal statement in [32] and [36]).
@@
Thus, for the logistic loss, the regime we are in is no longer just dependent on $\alpha$, the initialization scale, but rather on the combination of scale and training loss, which evolves in time. For each optimization accuracy, we can ask at what scale we would enter the kernel regime. The transition occurs at $\epsilon\sim\exp{(-\alpha^{2})}$. On the boundary, (under several strong assumptions), the behavior follows the $Q_{\alpha}$ function with parameter $\frac{\alpha}{\sqrt{\log(1/\epsilon)}}\,.$ Thus, the transition depends on the ordering of $\epsilon,\alpha$ limits. While it is true that we have an asymptotic result, it only kicks in when $\epsilon=10^{-1700}\,,$ so we do not get $\ell_{1}$ regularization in practice.
@@
Figure 20: This figure depicts how the predictor behaves with finite but not large initialization scale ($\alpha=2$). We know that in the long run, we will reach the minimum $\ell_{1}$ norm solution. When the optimization accuracy is finite, we will traverse the whole $Q_{\alpha}$ path. The dashed black line is the path of minimum $Q_{\alpha}$ margin solution: solution with margin 1 that minimizes $Q_{\alpha}$ for corresponding $\alpha$. As $\alpha$ increases, it seems we converge to this path. (Note that we don’t know how to prove this but it seems to hold empirically.) Importantly, while it is true asymptotically in $\epsilon\,$ that minimizing the regularizer in function space corresponds to minimizing the norm of the weights, the value of $\epsilon$ at which it starts to hold is completely impractical.
@@
Other control parameters (aside from initialization scale) include how early we stop training, the shape of the initialization (i.e., relative scale of the parameters), the step size, and the stochasticity. The latter has been studied through the lens of batch size and label noise.
@@ 6.4 Matrix Factorization Setting and Commutativity
We are still going to look at a linear model in $\beta$, but now $\beta$ is going to be a matrix. This is a standard least squares objective in $\beta$. The main difference is that now, instead of factorizing $\beta$ element-wise, we are going to employ the factorization $\beta=UV^{T}$. This formulation captures many things: matrix completion, matrix reconstruction from linear measurements, and even multi-task learning. In what follows, we will denote $\hat{L}$ the empirical loss that is being minimized.
@@
Observe that this is a linear transformation of the gradient. We would be interested in writing the transformation in terms of just $\beta$, and this is not possible here, as it turns out. Instead, we introduce an augmented variable:
@@
The local geometry of the space in which we search is given by left and right multiplication by $\tilde{\beta}\,.$ Now, we want to see if we can identify this as a Hessian map. Do the dynamics stay in a low-dimensional manifold?
As before, where we are depends on the integral of the residual so far. If matrices commute, we can solve the differential equation explicitly with a matrix exponential. The metric tensor is a Hessian map, which directly implies that the dynamcics occur in a low-dimension manifold. In this case, the result is very robust: it doesn’t depend on the residuals, nor the loss under which the residuals are computed. It is not robust to step size: a large step along the tangent space could lead to exiting the manifold now that it is curved. If the problem is non-commutative, the order in which we multiply on the left and right matters. $\beta(t)\,$ is then a “time-ordered exponential,” and we cannot ignore the ordering of the residuals. Even with just two data points, we can navigate the entire space instead of a low-dimension manifold. An analogy is parallel parking, where with the forward/backward and left/right controls we can somehow move straight to the right.
@@
7 Perspectives
We have seen three notions of implicit bias : one related to statistical guarantees of a model, benign overfitting and finally implicit bias related to optimization. The statistical aspect of implicit bias is well understood and falls within the framework of statistical learning theory. Leaving benign overfitting aside and focusing on optimization, we have demonstrated a general approach allowing to quantitatively study the implicit bias of a given descent algorithm for a chosen architecture. In particular, by identifying the potential that is implicitly minimized for a given problem, we are able to determine conditions to obtain an implicit bias akin to a kernel method or a model that is able to do feature selection or feature learning. While this approach is informative for the simple models we have considered, determining whether or not they generalize to more realistic neural network architectures remains unclear. Furthermore, it is possible that other forms of implicit bias exist, and that they are ultimately repsonsible for the empirical success of deep neural networks. In other words, we do not know if the implicit minimization of an effective potential is sufficient to explain most cases.
@@ Acknowledgements
These are notes from the lecture of Nathan Srebro given at the summer school “Statistical Physics & Machine Learning ”, that took place in Les Houches School of Physics in France from 4th to 29th July 2022. The school was organized by Florent Krzakala and Lenka Zdeborová from EPFL. Recordings of the lectures are available on the school’s website https://leshouches2022.github.io/.